OpenStudy (anonymous):

The chess club has 18 students. How many different teams of 4 can be formed for competition? 3,060 18,360 73,440

OpenStudy (anonymous):

\[^{18}C_4 ??\]

OpenStudy (anonymous):

Hey, wait, this is not answer, I am just checking if it is true as I am not sure about it.. :)

OpenStudy (anonymous):

hmmm dont know how to enter that into my calculator

OpenStudy (anonymous):

18 nCr 4, if you're using a TI-eighty something

OpenStudy (anonymous):

Or use Wolfram: http://www.wolframalpha.com/input/?i=18+choose+4

OpenStudy (anonymous):

not the answer tho how do i get it xD

OpenStudy (anonymous):

Calculate the value of: 18C4..

OpenStudy (anonymous):

how like what would that b on a calculator?

OpenStudy (anonymous):

\[\frac{18 \times 17 \times 16 \times 15}{4 \times 3 \times 2}\]

OpenStudy (anonymous):

\[\implies 180 \times 17 = ??\]

OpenStudy (anonymous):

3060 (": thanks

OpenStudy (anonymous):

It is okay.. ;)